# CVBench: Evaluating Cross-Video Synergies for Complex Multimodal Understanding and Reasoning

### v2

@@ A.2 QA Generation
We propose a hierarchical framework for generating cross-video question-answer pairs with three key stages for comprehensive evaluation. In the first stage, adaptive video segmentation is combined with GPT-4o to generate fine-grained spatiotemporal annotations, capturing critical elements such as object interactions and action sequences. The prompt template for generating cohesive video summaries is shown in Figure 6. In the second stage, a relationship extraction module constructs a cross-video association graph, identifying multidimensional relationships such as object evolution and event continuity. The cross-video relation extraction prompt is depicted in Figure 7. Based on 15 professional templates, the system generates diverse questions, ranging from basic matching to advanced reasoning, ensuring quality through multi-model validation and difficulty calibration. The QA pair prompt template is shown in Figure 8.
Figure 6: Prompt template for generating cohesive video summaries with GPT-4o. generate a cohesive summary of the following video by merging the segment-wise captions into a comprehensive temporal description that fully encapsulates the video’s content.
This framework integrates visual content analysis with logical reasoning, featuring a multi-layered quality control system that includes multi-model consensus validation and hallucination detection to ensure the verifiability and distinctiveness of the generated questions. Ultimately, the framework provides a comprehensive evaluation spectrum, ranging from single-hop to commonsense reasoning, with multidimensional metrics such as accuracy and reasoning depth. Figures 6, 7, and 8 present standardized prompt templates, ensuring the systematization and reproducibility of the evaluation process.
@@
Figure 7: Cross-video relation extraction prompt. Analyze the interconnections between multiple videos by combining their captions.
@@
Figure 8: QA pair prompt. Create quiz questions based on the interconnections between video captions.
A.3 More CVBench Examples Visualization
@@
CVBench’s 15 carefully designed subcategories of videos and questions evaluate key dimensions such as spatiotemporal correlation, entity-event association, and causal-logical reasoning. These tasks cover various challenges, from tracking object consistency across cameras to inferring causal relationships between multiple videos, detecting spatial annotation conflicts, and performing cross-modal data fusion. Each subcategory presents complex tasks that closely resemble real-world scenarios. More examples of these tasks are provided in Figure 9.

### v3

@@ Appendix B QA generation
We propose a hierarchical framework for generating cross-video question-answer pairs with three key stages for comprehensive evaluation. In the first stage, adaptive video segmentation is combined with GPT-4o to generate fine-grained spatiotemporal annotations, capturing critical elements such as object interactions and action sequences. The prompt template for generating cohesive video summaries is shown in Table 8. In the second stage, a relationship extraction module constructs a cross-video association graph, identifying multidimensional relationships such as object evolution and event continuity. The cross-video relation extraction prompt is depicted in Table 9. Based on 15 professional templates, the system generates diverse questions, ranging from basic matching to advanced reasoning, ensuring quality through multi-model validation and difficulty calibration. The QA pair prompt template is shown in Table 10.
Table 8: Prompt template for generating cohesive video summaries with GPT-4o.
You are an expert in analyzing videos. You will receive the video consisting of a series of frames. Your task is to carefully observe the details of the video frames and generate a detailed and accurate description of a video in English based on your observations. Instructions for writing the detailed description:
a). Focus on the key visual details such as appearance, motion, scene, actions, objects involved, and interactions between elements in the video.
b). Pay attention to the type of video and the important points like the order of events, appearance and actions of people or objects, and any significant changes, events or movements.
c). Note every detail in the video and give a clear and easy to understand description.
d). Indicators like ([Video number], Timestamp: [start time]-[end time]/[total time]), for example, (Video 1, Timestamp: 00:00-00:10/01:20), mean that the video is divided into parts, and the timestamps represent the start time and end time of the part. Please note that the change in Video number means that a new video has been inputed, the description of the new video has to be restarted, and the flow between the different parts of the same video has to ensure narrative continuity and avoid abrupt breaks in the description.
e). There is no need for markdown titles or emphasis such as ”**XXX**” or words such as ”(Frame X)” or ”(Video X, Timestamp: 00:XX-00:XX/0X:XX)”, emphasize the time in a uniform way, using a simple (00:XX-00:$X$) will do!
@@
Table 9: Cross-video relation extraction prompt. Analyze the interconnections between multiple videos by combining their captions.
You are an expert in analyzing interconnections of multiple videos. We extracted detailed captions (pertimestamp) for each video. For example, ’Video 1 : caption: 00:00-00:10: corresponds to the caption generated for the first video between 00:00 and 00:10.
Your task is to generate a detailed and accurate summary of interconnections between the videos (Video 1, Video2, etc.) as a paragraph, based on all the given video captions and videos. Make sure not to lose any important information. Use the following details to create a clear and complete narrative: Instructions for writing the detailed summary:
a). Pay strong attention to the interconnections between the videos, including differences, correlations, matches, etc. For instance, the correlations can be that the videos are split by a complete video, just in a different order; The differences can be different types of videos, or the same objects in them doing different actions; Matches can be the same person appearing in different videos. Check if events across videos follow a logical sequence, identify shared locations or objects across videos and note complementary angles.
b). Identify the theme of each video and use your intelligence to learn to infer hidden links.
c). Note the commonalities and contradictions between videos. The timestamps are just more detailed to allow you to understand the complete video, the focus is between the videos and not between the timestamps.
d). Do not mention that the information comes from captions.
e). Also, use your common sense to judge interconnections between the videos.
@@
Table 10: QA pair prompt. Create quiz questions based on the interconnections between video captions.
I would like you to act as a quizmaster who designs questions based on provided captions of each videos and summary of the interconnections between videos that should be grounded in real-life contexts and made more relevant to everyday realities. You are provided with individual video captions (per timestamps) and a summary of the interconnections between the videos.
You are to generate relevant QA pairs by analyzing key interconnections between videos while also referring to the detailed video summary. Your questions should specifically target the interconnections between the videos. Your question should be answerable, and checkable.Guidelines for Response:
1. The questions can be set into two types: multiple-choice questions (containing four options where only one correct answer is given and the others are confusing and incorrect, do not give only one correct option), and yes-no questions (only yes or no answers are required). If descriptions are ambiguous, (e.g., timeline or content), prioritize answer uniqueness. Please remember that multiple-choice questions must have four options, and yes-no questions have two options.
2. Avoid excessive similarity between options, and be more careful to avoid answers that can be inferred directly from the options.
3. When involving multiple subjects, use subject attributes (e.g.,“adult in red”) for reference. Avoid abstract identifiers like “subject 1” or “person A”.
4. Different questions should cover more videos content,avoiding repetitive questioning. 4. Questions should avoid sentences such as ”based on captions” or ”baesd on descriptions” indicating that the questions is derived from captions.
5. Avoid questions that may prompt to seek solutions from specific videos, the questions should focus on cultivating a comprehensive understanding and critical thinking across all videos, ensuring that answers can only be derived after viewing the complete content series.
6. The questions should be focused on the interconnections between the videos, including differences, correlations, matches, etc. Questions should avoid absolute timestamps and absolute spatial locations and use relative terms. For example, after drinking water, next to the TV, etc.
7. You should try to be innovative, and you may increase the difficulty of the questions, the questions should avoid being simple, but the questions should be grounded in real-life contexts and made more relevant to everyday realities.
This framework integrates visual content analysis with logical reasoning, featuring a multi-layered quality control system that includes multi-model consensus validation and hallucination detection to ensure the verifiability and distinctiveness of the generated questions. Ultimately, the framework provides a comprehensive evaluation spectrum, ranging from single-hop to commonsense reasoning, with multidimensional metrics such as accuracy and reasoning depth. Table 8, 9, and 10 present standardized prompt templates, ensuring the systematization and reproducibility of the evaluation process.
Appendix C More CVBench examples
@@
CVBench’s 15 carefully designed subcategories of videos and questions evaluate key dimensions such as spatiotemporal correlation, entity-event association, and causal-logical reasoning. These tasks cover various challenges, from tracking object consistency across cameras to inferring causal relationships between multiple videos, detecting spatial annotation conflicts, and performing cross-modal data fusion. Each subcategory presents complex tasks that closely resemble real-world scenarios. More examples of these tasks are provided in Fig. 6.